3x-4y=2
-2x+6y=-18

Answer by checkley79(3341)

3x-4y=2 multiply by 2
-2x+6y=-18 multiply by 3 & add.
6x-8y=4
-6x+18y=-54
-------------
10y=-50
y=-50/10
y=-5 ans.
3x-4*-5=2
3x+20=2
3x=2-20
3x=-18
x=-18/3
x=-6
Proof:
-2*-6+6*-5=-18
12-30=-18
-18=-18